# Training and Testing with Multiple Splits: A Central Limit Theorem for Split-Sample Estimators

### v1

@@ Abstract
As predictive algorithms grow in popularity, using the same dataset to both train and test a new model has become routine across research, policy, and industry. Sample-splitting attains valid inference on model properties by using separate subsamples to estimate the model and to evaluate it. However, this approach has two drawbacks, since each task uses only part of the data, and different splits can lead to widely different estimates. Averaging across multiple splits, I develop an inference approach that uses more data for training, uses the entire sample for testing, and improves reproducibility. I address the statistical dependence from reusing observations across splits by proving a new central limit theorem for a large class of split-sample estimators under arguably mild and general conditions. Importantly, I make no restrictions on model complexity or convergence rates. I show that confidence intervals based on the normal approximation are valid for many applications, but may undercover in important cases of interest, such as comparing the performance between two models. I develop a new inference approach for such cases, explicitly accounting for the dependence across splits. Moreover, I provide a measure of reproducibility for p-values obtained from split-sample estimators. Finally, I apply my results to two important problems in development and public economics: predicting poverty and learning heterogeneous treatment effects in randomized experiments. I show that my inference approach with repeated cross-fitting achieves better power than previous alternatives, often enough to find statistical significance that would otherwise be missed.
@@ 1 Introduction
As predictive algorithms become increasingly popular, using the same dataset to train and evaluate a new model has become routine across research, policy, and industry. Applications include estimating heterogeneous treatment effects from policy interventions, predicting future and contemporaneous outcomes, and building models for targeting. For example, the City of Chicago uses a machine learning (ML) algorithm to prioritize restaurant inspections most likely to find critical violations, while industry applications include personalized marketing ads and dynamic pricing. Despite their great potential, such procedures carry risks: algorithms may have low precision or negatively impact disadvantaged groups, for example by disproportionately inspecting restaurants owned by ethnic minorities. Thus, when training a new model, it is essential to evaluate its accuracy and distributional impacts before deploying it to perform or assist decision-making.
I study a setting in which an analyst (a researcher, policymaker, or industry practitioner) wishes to use the same dataset to both:
@@
In the restaurant inspections example, step (i) consists of training the ML model to predict whether an establishment has a critical violation, and a parameter of interest in step (ii) is the rate of correct classifications or the mean squared error. My setting imposes no complexity or convergence rate restrictions, thus accommodating popular machine learning algorithms such as random forests and neural networks. Specifically, I consider using multiple splits of the sample for conducting both tasks, derive a central limit theorem (CLT) for split-sample estimators, and provide valid confidence intervals for the parameters of interest in step (ii).
@@
Using the same observations for both steps (i) and (ii) creates a form of statistical dependence that makes inference challenging. This difficulty is often overcome by randomly splitting the sample into two, one part to train the model (training sample), and the other to evaluate its properties (evaluation sample). Since each task is conducted with separate data, such statistical dependence is not generated, and one can use standard approaches to inference. This procedure, however, has three drawbacks: it uses only part of the data for training the model, only part of the data for evaluating its properties, and different random splits can lead to widely different estimates and potentially affect statistical significance.
I study inference when averaging estimates across multiple sample splits, improving upon a regular 50/50 split by using more data for training, twice as much data for evaluation, and leading to better reproducibility properties. In empirical applications and Monte Carlo experiments, I show that these improvements often enable statistically significant conclusions that would otherwise be inconclusive. The main challenge of using multiple splits, for example with cross-fitting or repeated sample-splitting, is a new form of statistical dependence that complicates inference, since observations are used in both training and evaluation roles across different splits. I address this challenge by proving a new CLT for a large class of split-sample estimators, from which I develop new approaches to inference and a new reproducibility measure.
@@
To illustrate the technical challenges and empirical implications of my results, consider as a simple running example one of my applications, the problem of predicting poverty. Accurate out-of-sample poverty prediction is central to Development Economics for understanding poverty dynamics and designing targeted policy interventions. In this setting, I focus on assessing predictive accuracy as the natural starting point, though my framework applies more broadly. Consider a sample $D=(Y_{i},X_{i})_{i=1}^{n}$ of $n$ households, where $X$ are covariates and $Y$ is an indicator for being below the poverty line measured 13 years after the covariates. The goal is to use the sample to (i) train a model ${\hat{\eta}}(x)$ to predict poverty by estimating $P(Y=1|X=x)$, for example using a machine learning algorithm, and (ii) evaluate its accuracy, for example by estimating and calculating a confidence interval (CI) for the out-of-sample mean squared error (MSE)
@@
where $(Y_{new},X_{new})$ are out-of-sample observations from the same population as the sample. An alternative interpretation to $\theta_{\hat{\eta}}$ is given as follows. Suppose a policymaker is given a mandate to use the data to train a new model $^{\hat{\eta}}$, then apply it to predict $Y$ for the entire population. $\theta_{\hat{\eta}}$ reflects the accuracy that the trained model will have across the population. Note that $\theta_{\hat{\eta}}$ is data-dependent, and is thus different from targeting a parameter $\theta_{\eta_{0}}$ for some fixed $\eta_{0}$}. In the policy prediction example, the researcher is not interested in the out-of-sample accuracy of an ideal but unknown model $\eta_{0}$}. Instead, they are interested in the accuracy of the actually estimated model $^{\hat{\eta}}$.
In this context, CIs are often constructed using sample-splitting. If the entire sample is used for both tasks, standard CLTs do not apply to the average
@@
since the summands are not independent. For example, $Y_{1}-{\hat{\eta}}(X_{1})$ and $Y_{2}-{\hat{\eta}}(X_{2})$ are dependent since $^{\hat{\eta}}$ is estimated with both $(Y_{1},X_{1})$ and $(Y_{2},X_{2})$. A standard approach to handle this dependence is to impose complexity restrictions on how $^{\hat{\eta}}$ is estimated, such as Donsker conditions. These restrictions hold for simple procedures like ordinary least squares, but fail for complex machine learning algorithms frequently used in applied problems (chernozhukov2018double). Sample-splitting avoids this dependence without strong assumptions: randomly split $\{1,\dots,n\}$ into sets $\mathsf{s}_{1}$ and $\mathsf{s}_{2}$ of size for example $n/2$, use data in $\mathsf{s}_{1}$ to estimate ${\hat{\eta}}_{1}$ and data in $\mathsf{s}_{2}$ to calculate the average
@@
I make three main contributions. First, I prove a new central limit theorem for a large class of split-sample estimators under mild conditions. Specifically, I make no restrictions on the complexity of the models $^{\hat{\eta}}$, or on their rates of convergence or algorithmic stability. For sample-average estimators, my CLT follows under a standard moments condition and assuming that $^{\hat{\eta}}$ converges to an arbitrary limit, at any rate. I show that the normal approximation yields a valid CI in many applications, but may fail to do so in important cases of interest, such as comparing the performance between two models or some instances when $^{\hat{\eta}}$ converges to zero. My second contribution builds on the CLT to develop a new inference approach that covers such cases, explicitly accounting for the dependence across splits. I focus on the case of comparing the performance between two models, and discuss how the arguments apply more broadly to other cases. Finally, I develop a reproducibility measure for p-values obtained from split-sample estimators. It addresses a common concern: another researcher using the same dataset, but different splits, may reach a different conclusion about statistical significance. For a given (large) number of repetitions of sample-splitting/cross-fitting, my measure quantifies p-value reproducibility, assessing whether the number of repetitions is sufficiently large to ensure reproducible inference.
@@
I apply my inference approaches to two important problems in development and public economics: predicting poverty and learning heterogeneous treatment effects in randomized experiments. In the first application, using a panel from Ghana (ghanapaneldataset) and Monte Carlo experiments, I show that repeated cross-fitting outperforms previous alternative approaches in detecting predictive power for being below the poverty line 13 years ahead. In the second application, I revisit the experiment of karlan2007does on charitable giving and conduct Monte Carlo simulations, showing in both cases that my ensemble method achieves improved power for detecting heterogeneous treatment effects compared to previous alternatives.
@@ 1.1 Related Work
I join a large literature on inference using multiple splits of the data. The literature on risk estimation via cross-validation provides related results establishing asymptotic normality for sample-average estimators based on multiple splits. Like my approach, these CLTs target data-dependent parameters, but rely on different types of assumptions, and focus on the specific case of sample-averages. dudoit2005asymptotics consider estimators that average over all possible splits or cross-splits of the sample, assume bounded loss function and require $^{\hat{\eta}}$ to be loss consistent for a risk minimizing function, whereas I assume that $^{\hat{\eta}}$ converges to any limit. austern2020asymptotics; bayle2020cross provide CLTs under rate assumptions on the algorithmic stability of $^{\hat{\eta}}$. bayle2020cross provides two CLTs using estimators based on a single repetition of cross-fitting, one relying on rate condition for algorithmic stability, and the second requires a “conditional variance convergence” assumption that they verify using rates for loss stability. My result does not require verifying loss stability conditions, which may not be satisfied in some high-dimensional settings (bates2024cross), and my result allows for any ML algorithm as long as $^{\hat{\eta}}$ converges to any limit at any rate, in the sense established in Section 3. ledell2015computationally provides a CLT for the particular case of estimating the area under the curve (AUC) measure via cross-validation.
A different class of related results show asymptotic normality using cross-fitting when targeting parameters that are not data-dependent. These approaches require stronger conditions on $^{\hat{\eta}}$ that may not hold in general for nonparametric models with more than a handful of covariates, such as requiring $^{\hat{\eta}}$ to converge in probability at some specified rate (luedtke2016statistical; belloni2017program; chernozhukov2018double; benkeser2020improved; imai2025statistical). Theorem 3.1 requires no complexity restrictions, and assumes $^{\hat{\eta}}$ converges in probability to any limit at any rate.
@@ 2 Setup
A related estimand, also covered by my framework, is the difference in MSE between two groups, which is a measure of fairness (see, e.g., auerbach2024testing). Let $W=(Y,X,G)$, where $G\in\left\{a,b\right\}$ indicates group membership (e.g., racial groups). Then,
@@ Appendix H Additional Results
In the context of E.1, let $T=\left\{t\right\}$ and $\mathcal{P}=\left\{P\right\}$ be singletons, and let D.1 hold for $P$ and $f_{\eta,t}$ measurable with $f_{\eta,t}(w)<\infty$ for all $w$. Then, E.1 holds. ∎
@@
E.1(i) and E.1(ii) hold trivially. E.1(iii) holds by taking $f_{\eta,t}$ as its own envelope. The uniform integrability condition E.1(iv) is implied by the $2+\delta$ assumption D.1(i). E.1(v) holds trivially. E.1(vi) holds since both covering and bracketing numbers are equal to $1$ with singleton $T$. Finally, E.2 follows since
@@
as established under D.1(ii) in the proof of Theorem D.1 D.13, since convergence in $L_{1}$ implies convergence in probability. ∎

### v2

@@ Abstract
As predictive algorithms grow in popularity, using the same dataset to both train and test a new model has become routine across research, policy, and industry. Sample-splitting attains valid inference on model properties by using separate subsamples to estimate the model and to evaluate it. However, this approach has two drawbacks, since each task uses only part of the data, and different splits can lead to widely different estimates. Averaging across multiple splits, I develop an inference approach that uses more data for training, uses the entire sample for testing, and improves reproducibility. I address the statistical dependence from reusing observations across splits by proving a new central limit theorem for a large class of split-sample estimators under arguably mild and general conditions. Importantly, I make no restrictions on model complexity or convergence rates. I show that confidence intervals based on the normal approximation are valid for many applications, but may undercover in important cases of interest, such as comparing the performance between two models. I develop a new inference approach for such cases, explicitly accounting for the dependence across splits. Moreover, I provide a measure of reproducibility for p-values obtained from split-sample estimators. Finally, I apply my results to two important problems in development and public economics: predicting poverty and learning heterogeneous treatment effects in randomized experiments. I show that my inference approach with repeated cross-fitting achieves better power than existing alternatives, often enough to reveal statistical significance that would otherwise be missed.
@@ 1 Introduction
As predictive algorithms transform empirical economics, policy, and industry, it is now routine to use the same dataset to train and evaluate a new model. For example, when training a machine learning algorithm to predict treatment effects, create a targeted policy rule, or automate consumer credit scoring, it is essential to evaluate the quality of the predictions and assess whether implementation would generate disparate impact across demographic groups. However, standard methods for training and evaluating typically waste part of the data by splitting the sample into training and testing sets. I develop a new inference approach that uses the entire sample for both tasks by combining multiple splits, improving statistical power and reproducibility. I provide valid confidence intervals under weak conditions for model properties such as accuracy and fairness calculated using cross-fitting, repeated sample-splitting or repeated cross-fitting.
Specifically, I study a setting in which an analyst (a researcher, policymaker, or industry practitioner) wishes to use the same dataset to both:
@@
For example, consider a government using machine learning (ML) to target recipients of a poverty alleviation program. Step (i) consists of training a model to predict, for example, families at higher risk of falling below the poverty line, while step (ii) consists of constructing a confidence interval for the out-of-sample mean squared error or rate of correct classifications.
@@
Using the same observations for both steps (i) and (ii) creates a form of statistical dependence that makes inference challenging. For example, standard central limit theorems (CLTs) assume independence, which is violated in this setting. This difficulty is often overcome by randomly splitting the sample into two, one part to train the model (training sample), and the other to evaluate its properties (evaluation sample). Since each task is conducted with separate data, such statistical dependence is not generated, and one can use standard approaches to inference. This procedure, however, has three drawbacks: it uses only part of the data for training the model, only part of the data for evaluating its properties, and different random splits can lead to widely different estimates and potentially affect statistical significance.
My main contribution is an inference approach that averages estimates across multiple sample splits, improving upon a standard 50/50 split by using more data for training, using twice as much data for evaluation, and improving reproducibility. In empirical applications and Monte Carlo experiments, I show that these improvements often reveal statistical significance that would otherwise be missed. The main challenge of using multiple splits is a new form of statistical dependence due to reusing observations in both training and evaluation roles across different splits. I address this challenge by proving a new CLT for a large class of split-sample estimators under weak conditions. My CLT builds on previous literature in two key dimensions: (i) it applies to a large class of estimators and split-sample procedures, and (ii) it imposes no restrictions on model complexity or rates of convergence and stability, only requiring that the estimated model converges to an arbitrary limit at any rate. This generality is crucial for accommodating popular ML algorithms, such as random forests or neural networks. Moreover, I characterize when confidence intervals based on the normal approximation are valid, showing they may fail in important cases such as comparing the performance of two models or learning features of heterogeneous treatment effects. I develop a new inference approach for such cases that explicitly accounts for the dependence across splits, leveraging my CLT. Finally, I develop a reproducibility measure for p-values from split-sample procedures, quantifying whether the number of repetitions is sufficiently large to ensure reproducible inference.
@@
To illustrate the technical challenges and empirical implications of my results, consider the problem of predicting poverty as a simple running example, which is one of my applications. Accurate out-of-sample poverty prediction is central to Development Economics for understanding poverty dynamics and designing targeted interventions. I focus on assessing predictive accuracy as the natural starting point, though my framework applies more broadly. Consider a sample $D=(Y_{i},X_{i})_{i=1}^{n}$ of $n$ households, where $X_{i}$ are covariates and $Y_{i}$ is a binary indicator equal to one if household $i$ is below the poverty line 13 years after the covariates were measured. The goal is to use the sample to (i) train a model ${\hat{\eta}}(x)$ to predict poverty by estimating $P(Y=1|X=x)$, for example using a machine learning algorithm, and (ii) evaluate its accuracy, for example by estimating and calculating a confidence interval (CI) for the out-of-sample mean squared error (MSE)
@@
where $(Y_{new},X_{new})$ is an out-of-sample observation drawn from the same distribution as the sample. Note that $\theta_{\hat{\eta}}$ is data-dependent, and is thus different from targeting a parameter $\theta_{\eta_{0}}$ for some fixed $\eta_{0}$}. In the policy prediction example, the researcher is not interested in the out-of-sample accuracy of an ideal but unknown model $\eta_{0}$}. Instead, they are interested in the accuracy of the actually estimated model $^{\hat{\eta}}$.
In this context, confidence intervals are often constructed using sample-splitting. If the entire sample is used for both tasks, standard CLTs do not apply to the average
@@
since the summands are not independent. For example, $Y_{1}-{\hat{\eta}}(X_{1})$ and $Y_{2}-{\hat{\eta}}(X_{2})$ are dependent since $^{\hat{\eta}}$ is estimated with both $(Y_{1},X_{1})$ and $(Y_{2},X_{2})$. A standard approach to handle this dependence is to impose complexity restrictions on how $^{\hat{\eta}}$ is estimated, such as Donsker conditions. These restrictions hold for simple procedures like ordinary least squares, but fail for complex machine learning algorithms frequently used in applied problems (chernozhukov2018double). Sample-splitting avoids this dependence without strong assumptions: randomly split $\{1,\dots,n\}$ into sets $\mathsf{s}_{1}$ and $\mathsf{s}_{2}$ of size for example $n/2$, use data in $\mathsf{s}_{1}$ to estimate ${\hat{\eta}}_{1}$, and data in $\mathsf{s}_{2}$ to calculate the average
@@
I make three main contributions. First, I prove a new central limit theorem for a large class of split-sample estimators under mild conditions. Specifically, I make no restrictions on the complexity of the models $^{\hat{\eta}}$, or on their rates of convergence or algorithmic stability. For sample-average estimators, my CLT follows under a standard moments condition and assuming that $^{\hat{\eta}}$ converges to an arbitrary limit, at any rate. I show that the normal approximation yields a valid CI in many applications, but may fail to do so in important cases of interest, such as comparing the performance between two models or some instances when $^{\hat{\eta}}$ converges to a constant. My second contribution builds on the CLT to develop a new inference approach that covers such cases, explicitly accounting for the dependence across splits. I focus on the case of comparing the performance between two models, and discuss how the arguments apply more broadly to other cases. Finally, I develop a reproducibility measure for p-values obtained from split-sample estimators. It addresses a common concern: another researcher using the same dataset, but different splits, may reach a different conclusion about statistical significance. For a given (large) number of repetitions of sample-splitting/cross-fitting, my measure quantifies p-value reproducibility, assessing whether the number of repetitions is sufficiently large to ensure reproducible inference.
@@
I apply my inference approaches to two important problems in development and public economics: predicting poverty and learning heterogeneous treatment effects in randomized experiments. In the first application, using a panel from Ghana (ghanapaneldataset) and Monte Carlo experiments, I show that repeated cross-fitting outperforms previous alternative approaches in detecting predictive power for being below the poverty line 13 years ahead. In the second application, I revisit the experiment of karlan2007does on charitable giving and conduct Monte Carlo simulations, and show that my ensemble method achieves improved power for detecting heterogeneous treatment effects compared to previous alternatives.
@@ 1.1 Related Work
I contribute to the literature on inference using multiple splits of the data. The literature on risk estimation via cross-validation provides related results establishing asymptotic normality for sample-average estimators based on multiple splits. Like my approach, these CLTs target data-dependent parameters, but rely on different types of assumptions and focus on the particular case of sample-averages. dudoit2005asymptotics consider estimators that average over all possible splits or cross-splits of the sample, assume bounded loss function and require $^{\hat{\eta}}$ to be loss consistent for a risk minimizing function, whereas I assume that $^{\hat{\eta}}$ converges to an arbitrary limit. austern2020asymptotics and bayle2020cross provide CLTs under rate assumptions on the algorithmic stability of $^{\hat{\eta}}$. bayle2020cross provide two CLTs using estimators based on a single repetition of cross-fitting, one relying on rate conditions for algorithmic stability, and the second requires a “conditional variance convergence” assumption that they verify using rates for loss stability. My result does not require verifying loss stability conditions, which may not be satisfied in some high-dimensional settings (bates2024cross), and my result allows for any ML algorithm as long as $^{\hat{\eta}}$ converges to an arbitrary limit at any rate, in the sense established in Section 3. ledell2015computationally provide a CLT for the particular case of estimating the area under the curve (AUC) measure via cross-validation, and andrews2022transfer derive confidence intervals for the different but related problem of learning the transfer error of a model across domains. Moreover, I document that asymptotic normality of split-sample estimators does not immediately lead to valid inference for the important problem of comparing the performance between two models, and I construct a new inference approach for this case that explicitly incorporates the dependence across splits.
A different class of related results show asymptotic normality using cross-fitting when targeting parameters that are not data-dependent. These approaches require stronger conditions on $^{\hat{\eta}}$ that may not hold in general for nonparametric models with more than a handful of covariates, such as requiring $^{\hat{\eta}}$ to converge in probability at some specified rate (luedtke2016statistical; belloni2017program; chernozhukov2018double; benkeser2020improved; imai2025statistical). Leveraging the data-dependent parameter of interest, my CLT (Theorem 3.1) requires no complexity restrictions, and assumes $^{\hat{\eta}}$ converges in probability to any limit at any rate.
@@ 2 Setup
A related estimand, also covered by my framework, is the difference in MSE between two groups, which is a measure of fairness (auerbach2024testing; liang2024algorithm; liu2024inference). Let $W=(Y,X,G)$, where $G\in\left\{a,b\right\}$ indicates group membership (e.g., racial groups). Then,
@@ Appendix H Additional Results
In the context of Assumption E.1, let $T=\left\{t\right\}$ and $\mathcal{P}=\left\{P\right\}$ be singletons, and let Assumption D.1 hold for $P$ and $f_{\eta,t}$ measurable with $f_{\eta,t}(w)<\infty$ for all $w$. Then, Assumption E.1 holds. ∎
@@
Assumption E.1(i) and Assumption E.1(ii) hold trivially. Assumption E.1(iii) holds by taking $f_{\eta,t}$ as its own envelope. The uniform integrability condition Assumption E.1(iv) is implied by the $2+\delta$ assumption Assumption D.1(i). Assumption E.1(v) holds trivially. Assumption E.1(vi) holds since both covering and bracketing numbers are equal to $1$ with singleton $T$. Finally, Assumption E.2 follows since
@@
as established under Assumption D.1(ii) in the proof of Theorem D.1 D.13, since convergence in $L_{1}$ implies convergence in probability. ∎